40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Jan 4 12:49:18 2015, Steve, Update, SUS, recent earthquakes recentEQ.png
    Reply  Mon Jan 5 12:42:09 2015, ericq, Update, SUS, BS misbehaving 
Message ID: 10852     Entry time: Mon Jan 5 12:42:09 2015     In reply to: 10850
Author: ericq 
Type: Update 
Category: SUS 
Subject: BS misbehaving 

The BS was showing some excess motion. I think I've fixed it. Order of operations:

  • The DC PIT bias from previous ASS runs was at ~500, I zeroed this and aligned the BS to be centered on its oplev QPD with DC alignment sliders
  • I squished the gold box cables. This changed the alignment slightly, and brought the UR voltage back to a normal value. Excess motion still existed
  • I found that the the C1:SUS-BS_LRSEN filter had HOLD OUTPUT enabled. I turned it off. All seems well. 

I'm not sure how this might have gotten switched on...

ELOG V3.1.3-